OpenStudy (anonymous):

What is the solution set for |3x + 12|+ 5 = 23? x = 2 and x = 10 x = –2 and x = 10 x = –2 and x = –10 x = 2 and x = –10

OpenStudy (anonymous):

1) subtract 5 and get \[|3x+12|=18\] 2) make two equations \[3x+12=-18\] and \[3x+12=18\] and solve those

OpenStudy (lopezking1):

|3x + 12|+ 5 = 23 |3x + 12| =18 3x+12 = 18, 3x + 12= -18 3x = 6, 3x= -30 x = 2, x= -10

OpenStudy (anonymous):

each takes exactly two steps, and they aer the same 1) subtract 12 2) divide by 3

OpenStudy (anonymous):

so D

OpenStudy (lopezking1):

Mhmm :)